Johnson Family

  • Married couple
  • 3 kids all under (16 year old twins, 21 year old college student who lives at home)
  • Wife earns $75,000/yr
  • Husband earns $24,000/yr
  • They own a home and paid $8,200 in mortgage interest last year
  • They paid $3,600 in real estate taxes
  • Family donated $1,500 to the church last year
  • Family Contributed $400 to Goodwill
  • Husband paid $2,700 in student loan interest last year
  • Family paid a combined $9,000 in federal income tax through their paychecks
  • Own 3 cars that are all paid off
  • Husband Contributed $6,000 to a traditional IRA last year
  • Wife Contributed $500 to a Traditional IRA

1. What is the Johnson Family AGI?

2. What would the standard deduction be for the Johnson family?

3. What is the total of itemized deductions the Johnson family can use?

4. What is the total for exemptions available for the Johnson family?

5. What is the Johnson family Taxable Income?

6. What is the Johnson family marginal tax rate?

7. How much will the Johnsons owe/get in refund for their tax return?

8. What is the average tax rate for the Johnson family?
Answer #1

PART-1

Adjusted Gross Income (AGI) is an individual's total gross income minus specified deductions.

AGI = Gross Income less (Itemized Deductions or Standard Deduction whichever is more)

GROSS INCOME

Salary earned by wife + salary earned by husband

i.e 75000+24000 = 99000 dollars

Therefore, Gross Income of Johnson's Family is 99000 Dollars.

ITEMIZED DEDUCTIONS

Home mortgage interest = 8200 Dollars

Real estate tax = 3200 Dollars (Max 10000 dollars)

Donation to church = 1500 Dollars

Goodwill donations = 400 Dollars

Student loan interest = 2500 Dollars (max limit)

Contribution to TRADITIONAL IRA = 6000+500

= 6500 Dollars(Max limit = 11000 dollars)

Therefore, Total itemized deduction = 8200 + 3200+1500+400+2500+6500

= 22300 Dollars

STANDARD DEDUCTION

Standard deduction for married couples filing jointly is 24400 .

Therefore, AGI = Gross Income - ( Itemized deduction or Standard deduction whichever is more)

= 99000 - (22300 or 24400 )

= 99000-24400

= 74600 Dollars

PART-2

Standard deduction amount will increase to 12200 for individuals , 18350 for head of household and 24400 for married couples filing jointly.

In this question, Johnson family is a married couple filing jointly. Hence, Standard Deduction = 24400 Dollars.

PART-3

Itemized deductions can only be used if the deductions available is more than the Standard deduction provided .

Here, the total of Itemized deduction is 22300 Dollars and Standard deduction available is 24400 Dollars.

Hence Johnson family cannot use Itemized deduction as the Standard deduction is more than Itemized deduction.

PART-4

As per the new Income tax rules personal exemption will be zero. Hence there will be no exemption available for Johnson family.